OpenStudy (anonymous):

Will fan if solved!!! Find theta to 4 significant digits for 0

OpenStudy (poopsiedoodle):

@Hero @nincompoop

OpenStudy (kmeis002):

Note that \[ cos(\theta ) = -.8759\] Is asking what angle will give you an x-coordinate of -.8759, so you have to use inverse trig function arccos (sometimes noted as cos^-1) Take the arccos of each side: \[ arccos(cos(\theta)) = arccos(-.8759) \] Since arccos is the inverse of cos, it will undo what cos has done so we get: \[ \theta = arccos(-.8795) = cos^{-1}(-.8795)\] Gotta be careful, since you have TWO potential angles that could result in your x coordinate of -.8795, see picture attached) Theta_1 is what the calculator will return, but from the image, you can see we can derive the second solution from the equation \[ \theta_2 = 2 \pi - \theta_1 \] I ran the numbers and obtained one of your solutions, so it should work. I hope I made it clear enough.

OpenStudy (anonymous):

ahhh thanks

OpenStudy (anonymous):

use arccos right?